OpenStudy (anonymous):

how do you know that 7/8 - 1/4 will be more than 1/2

OpenStudy (anonymous):

Calculate it: 1/2 = 4/8 7/8 - 1/4 = 7/8 -2/8 = 5/8 5/8 > 4/8, Therefore, 7/8 - 1/4 > 1/2

OpenStudy (anonymous):

7/8 - 2/8 = 5/8 1/2=4/8